The resultant of two forces at right angle is 17N if the maximum resultant is 23N the greater force is

Grade:

2 Answers

Arun
5 years ago
If F1 and F2 are two forces with resultant "R'' and ∅ being angle between them
           R=(√F1²+F2²+2F1F2cos∅)
If  F1 and F2 are at 90°
   R=(√F1²+F2²+2F1F2cos90)
      =(√F1²+F2²)
If R is maximum ⇒⇒cos∅ is maximum
Maximum value of cos∅= 1
So,
 R=(√F1²+F2²+2F1F2cos0)=(√F1²+F2²+2F1F2)
~~~~~~~~~~~~~~~~~~~~~~~~~
According to question 
Resultant of F1 and F2 at right angle is 17N
 17=(√F1²+F2²)
  289=(F1²+F2²)
~~~~~~~~~~~~~~~~~~~
maximum resultant is 23N  
23= (√F1²+F2²+2F1F2)
 529=(F1²+F2²+2F1F2)
 289=(F1²+F2²)
 529=289+2F1F2
 529-289=2F1F2
 240=2F1F2
 120=F1F2
Now 
(F1+F2)²=F1²+F2²+2F1F2
F1+F2=√289+240
F1+F2=23
Similarly (F1-F2)²=F1²+F2²-2F1F2
F1-F2=√289-240
F1-F2=√289-240=√49=7
F1-F2=7
F1+F2=23
---------------
2F1=30
F1=15 F2=8
So greater force is 15N
Regards
Arun (askIITians forum expert)
Eshan
5 years ago
Let the forces be\vec{F_1}and\vec{F_2}.
The resultant of two forces kept perpendicular is\sqrt{F_1^2+F_2^2}=17N
Maximum resultant is possible when the two forces at in the same direction.

Therefore,F_1+F_2=23N
HenceF_1=15N, F_2=8N